Klose Outfitters, Inc. believes that its optimal capital structure consists of 60% common equity and 40% debt, and its tax rate is 40%. Klose must raise additional capital to fund its upcoming expansion. The firm will have $2 million of retained earnings with a cost of 12%. New common stock in an amount up to $6 million would have a cost of 16%. Furthermore, Klose can raise up to $3 million of debt at an interest rate of 10% and an additional $4 million of debt at 11%. The CFO estimates that a proposed expansion would require an investment of $5.9 million. What is the WACC for the last dollar raised to complete the expansion?
Answer #1

Common equity required = common equity%*project cost = 0.6*5.9 = 3.54m

debt required = debt%*project cost = 0.4*5.9 =2.36 m

retained earnings of 2m not enough to meet equity requirement

therefore new common equity to raise @ 16%

debt required falls in first bucket of 3 m

therefore cost of debt = 10%

After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 10*(1-0.4)
= 6
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=6*0.4+16*0.6
WACC =12%